# EW D-optimal Designs for Experiments with Mixed Factors

### v1

@@ 3.1 Lift-one algorithm for EW D-optimal designs with discrete factors only
${\mathbf{X}}_{\mathbf{x}}=\begin{pmatrix}{\mathbf{h}}_{1}^{T}({\mathbf{x}})&\boldsymbol{0}^{T}&\cdots&\boldsymbol{0}^{T}&{\mathbf{h}}_{c}^{T}({\mathbf{x}})\\ \boldsymbol{0}^{T}&{\mathbf{h}}_{2}^{T}({\mathbf{x}})&\ddots&\vdots&\vdots\\ \vdots&\ddots&\ddots&\boldsymbol{0}^{T}&{\mathbf{h}}_{c}^{T}({\mathbf{x}})\\ \boldsymbol{0}^{T}&\cdots&\boldsymbol{0}^{T}&{\mathbf{h}}_{J-1}^{T}({\mathbf{x}})&{\mathbf{h}}_{c}^{T}({\mathbf{x}})\\ \boldsymbol{0}^{T}&\cdots&\cdots&\boldsymbol{0}^{T}&\boldsymbol{0}^{T}\\ \end{pmatrix}_{J\times p}$ (5)
@@
with pre-determined predictor functions ${\mathbf{h}}_{j}=(h_{j1},\ldots,h_{jp_{j}})^{T}$, $j=1,\ldots,J-1$, ${\mathbf{h}}_{c}=(h_{1},\ldots,h_{p_{c}})^{T}$, and ${\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\theta})=\left(u_{st}^{\mathbf{x}}(\boldsymbol{\theta})\right)_{s,t=1,\ldots,J}$ Then $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}={\mathbf{X}}_{\mathbf{x}}^{T}E\{{\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\mathbf{x}}$ and $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}={\mathbf{X}}_{\mathbf{x}}^{T}\hat{E}\{{\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\mathbf{x}}$Box
Theorem 3.1.
@@
For MLMs under Assumptions (A1) and (B3), suppose all the predictor functions ${\mathbf{h}}_{1},\ldots,{\mathbf{h}}_{J-1}$ and ${\mathbf{h}}_{c}$ defined in (5) are continuous with respect to all continuous factors of $\mathbf{x}\in{\cal X}$, and the parameter space $\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$ is bounded. Then there exists an EW D-optimal design $\boldsymbol{\xi}$ that contains no more than $p(p+1)/2$ support points, and the design obtained by Algorithm 1 must be EW D-optimal.
@@
The boundedness of $\boldsymbol{\Theta}$ in Theorem 3.1 is a practical requirement. For typical applications, the boundedness of a working parameter space is often needed for a numerical search for parameter estimates or a theoretical derivation for desired properties (Ferguson, 1996).
Example 3.2.
Generalized linear models (GLM):For a generalized linear model (McCullagh and Nelder, 1989; Dobson and Barnett, 2018) defined by
$E(Y_{i})=\mu_{i}\mbox{ and }\eta_{i}=g(\mu_{i})={\mathbf{X}}_{i}^{T}\boldsymbol{\theta}\ ,$
@@
where $g$ is a given link function, ${\mathbf{X}}_{i}={\mathbf{h}}({\mathbf{x}}_{i})=(h_{1}({\mathbf{x}}_{i}),\ldots,h_{p}({\mathbf{x}}_{i}))^{T}$, $i=1,\ldots,m$ (see, for example, Section 4 in Huang et al. (2024)). Then ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})=\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}\}\cdot{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}$, where $\nu=\{(g^{-1})^{\prime}\}^{2}/s$ with $s(\eta_{i})={\rm Var}(Y_{i})$ (see Table 5 in the Supplementary Material of Huang et al. (2025) for examples of $\nu$). To calculate $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ under GLMs, it is enough to replace $\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}\}$ with $E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]=\int_{\boldsymbol{\Theta}}\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}\}Q(d\boldsymbol{\theta})$ or $\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]$ $=B^{-1}\sum_{j=1}^{B}\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\hat{\boldsymbol{\theta}}_{j}\}$. That is, $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}=E\left[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}\right]\cdot{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}$, and $\hat{E}\{{\mathbf{F}}({\mathbf{x}},$ $\boldsymbol{\Theta})\}=\hat{E}\left[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}\right]\cdot{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}$.
@@
To calculate ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})$ under GLMs, instead of using $\sum_{i=1}^{m}w_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$, one may use its matrix form ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})={\mathbf{X}}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\theta}){\mathbf{X}}_{\boldsymbol{\xi}}$ , where ${\mathbf{X}}_{\boldsymbol{\xi}}=({\mathbf{h}}({\mathbf{x}}_{1}),\ldots,{\mathbf{h}}({\mathbf{x}}_{m}))^{T}\in\mathbb{R}^{m\times p}$, and ${\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\theta})={\rm diag}\{w_{1}\nu\{{\mathbf{h}}({\mathbf{x}}_{1})^{T}{\boldsymbol{\theta}}\},\ldots,$ $w_{m}\nu\{{\mathbf{h}}({\mathbf{x}}_{m})^{T}{\boldsymbol{\theta}}\}$. Then, under a GLM,
$E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})\}={\mathbf{X}}_{\boldsymbol{\xi}}^{T}E\{{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\boldsymbol{\xi}}\>\mbox{ and }\>\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})\}={\mathbf{X}}_{\boldsymbol{\xi}}^{T}\hat{E}\{{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\boldsymbol{\xi}}\ .$
Note that the matrix form is much more efficient especially when programming in R. $\Box$
Theorem 3.2.

### v2

@@ 3.2 EW ForLion algorithm and rounding algorithm
${\mathbf{X}}_{\mathbf{x}}=\begin{pmatrix}{\mathbf{h}}_{1}^{T}({\mathbf{x}})&\boldsymbol{0}^{T}&\cdots&\boldsymbol{0}^{T}&{\mathbf{h}}_{c}^{T}({\mathbf{x}})\\ \boldsymbol{0}^{T}&{\mathbf{h}}_{2}^{T}({\mathbf{x}})&\ddots&\vdots&\vdots\\ \vdots&\ddots&\ddots&\boldsymbol{0}^{T}&{\mathbf{h}}_{c}^{T}({\mathbf{x}})\\ \boldsymbol{0}^{T}&\cdots&\boldsymbol{0}^{T}&{\mathbf{h}}_{J-1}^{T}({\mathbf{x}})&{\mathbf{h}}_{c}^{T}({\mathbf{x}})\\ \boldsymbol{0}^{T}&\cdots&\cdots&\boldsymbol{0}^{T}&\boldsymbol{0}^{T}\\ \end{pmatrix}_{J\times p}$ (3.7)
@@
with predetermined predictor functions ${\mathbf{h}}_{j}=(h_{j1},\ldots,h_{jp_{j}})^{T}$, $j=1,\ldots,J-1$, ${\mathbf{h}}_{c}=(h_{1},\ldots,h_{p_{c}})^{T}$, and ${\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\theta})=\left(u_{st}^{\mathbf{x}}(\boldsymbol{\theta})\right)_{s,t=1,\ldots,J}$ Then $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}={\mathbf{X}}_{\mathbf{x}}^{T}E\{{\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\mathbf{x}}$ and $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}={\mathbf{X}}_{\mathbf{x}}^{T}\hat{E}\{{\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\mathbf{x}}$Box
Theorem 3.
@@
For MLMs under Assumptions (A1) and (B3), suppose all the predictor functions ${\mathbf{h}}_{1},\ldots,{\mathbf{h}}_{J-1}$ and ${\mathbf{h}}_{c}$ defined in (3.7) are continuous with respect to all continuous factors of $\mathbf{x}\in{\cal X}$, and the parameter space $\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$ is bounded. Then there exists an EW D-optimal design $\boldsymbol{\xi}$ that contains no more than $p(p+1)/2$ support points, and the design obtained by Algorithm 1 must be EW D-optimal.
@@
The boundedness of $\boldsymbol{\Theta}$ in Theorem 3 is a practical requirement. For typical applications, the boundedness of a working parameter space is often needed for numerical searches for parameter estimates or theoretical derivations for desired properties (Ferguson, 1996).
Example 4.
@@
Generalized linear models (GLM):A generalized linear model (McCullagh and Nelder, 1989; Dobson and Barnett, 2018) can be defined by $E(Y_{i})=\mu_{i}$ and $\eta_{i}=g(\mu_{i})={\mathbf{X}}_{i}^{T}\boldsymbol{\theta}$, where $g$ is a given link function, ${\mathbf{X}}_{i}={\mathbf{h}}({\mathbf{x}}_{i})=(h_{1}({\mathbf{x}}_{i}),\ldots,h_{p}({\mathbf{x}}_{i}))^{T}$, $i=1,\ldots,m$, and $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$ (see, for example, Section 4 in Huang et al. (2024)). Then ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})=\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}\}\cdot{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}$, where $\nu=\{(g^{-1})^{\prime}\}^{2}/s$ with $s(\eta_{i})={\rm Var}(Y_{i})$ (see Table 5 in the Supplementary Material of Huang et al. (2025) for examples of $\nu$). To calculate ${\mathbf{F}}_{\mathbf{x}}=E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ under GLMs, it is enough to replace $\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}\}$ with $E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]=\int_{\boldsymbol{\Theta}}\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\theta}\}Q(d\boldsymbol{\theta})$ or $\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]$ $=B^{-1}\sum_{j=1}^{B}\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\hat{\boldsymbol{\theta}}_{j}\}$. That is, $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}=E\left[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}\right]\cdot{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}$, and $\hat{E}\{{\mathbf{F}}({\mathbf{x}},$ $\boldsymbol{\Theta})\}=\hat{E}\left[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}\right]\cdot{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}})^{T}$.
@@
To calculate ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})$ under GLMs, instead of $\sum_{i=1}^{m}w_{i}{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\theta})$, we recommend its matrix form ${\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})={\mathbf{X}}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\theta}){\mathbf{X}}_{\boldsymbol{\xi}}$ , where ${\mathbf{X}}_{\boldsymbol{\xi}}=$ $({\mathbf{h}}({\mathbf{x}}_{1}),\ldots,$ ${\mathbf{h}}({\mathbf{x}}_{m}))^{T}\in\mathbb{R}^{m\times p}$, and ${\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\theta})={\rm diag}\{w_{1}\nu\{{\mathbf{h}}({\mathbf{x}}_{1})^{T}{\boldsymbol{\theta}}\},\ldots,w_{m}\nu\{{\mathbf{h}}({\mathbf{x}}_{m})^{T}{\boldsymbol{\theta}}\}\}$. Then $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})\}={\mathbf{X}}_{\boldsymbol{\xi}}^{T}E\{{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\Theta})\}$ ${\mathbf{X}}_{\boldsymbol{\xi}}$ and $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})\}={\mathbf{X}}_{\boldsymbol{\xi}}^{T}\hat{E}\{{\mathbf{W}}_{\boldsymbol{\xi}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{\boldsymbol{\xi}}$ The matrix form is much more efficient when programming in R. $\Box$
Theorem 4.
@@
3.4 First-order derivative of sensitivity function
To implement the EW ForLion algorithm (Algorithm 1) for EW D-optimal designs, if there is at least one continuous factor, we need to calculate the first-order derivative of the sensitivity function $d({\mathbf{x}},\boldsymbol{\xi})$ in Theorem 2. Following the simplified notation in Section 3.2, $d({\mathbf{x}},\boldsymbol{\xi})={\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})$.
Similarly to Theorem 3 in Huang et al. (2024), we obtain the following theorem for multinomial logistic models (MLM) under EW D-optimality.
Theorem 5.
For an MLM model under Assumptions (A1) and (B3), consider $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ satisfying $|{\mathbf{F}}(\boldsymbol{\xi})|>0$. If we denote the $p\times p$ matrix ${\mathbf{F}}(\boldsymbol{\xi})^{-1}=(\mathbf{E}_{st})_{s,t=1,\ldots,J}$ with submatrix $\mathbf{E}_{st}\in\mathbf{R}^{p_{s}\times p_{t}}$, and $p_{J}=p_{c}$ , ${\mathbf{h}}_{j}^{\mathbf{x}}={\mathbf{h}}_{j}({\mathbf{x}})$, ${\mathbf{h}}_{c}^{\mathbf{x}}={\mathbf{h}}_{c}({\mathbf{x}})$, $u_{st}^{\mathbf{x}}=E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}$ or $\hat{E}\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}$ for simplicity, then
$d(\mathbf{x},\boldsymbol{\xi})$ $=$ $\sum_{t=1}^{J-1}u_{tt}^{\mathbf{x}}\left(\mathbf{h}_{t}^{\mathbf{x}}\right)^{T}\mathbf{E}_{tt}\mathbf{h}_{t}^{\mathbf{x}}+\sum_{s=1}^{J-1}\sum_{t=1}^{J-1}u_{st}^{\mathbf{x}}\left(\mathbf{h}_{c}^{\mathbf{x}}\right)^{T}\mathbf{E}_{JJ}\mathbf{h}_{c}^{\mathbf{x}}$
$+$ $2\sum_{s=1}^{J-2}\sum_{t=s+1}^{J-1}u_{st}^{\mathbf{x}}\left(\mathbf{h}_{t}^{\mathbf{x}}\right)^{T}\mathbf{E}_{st}\mathbf{h}_{s}^{\mathbf{x}}+2\sum_{s=1}^{J-1}\sum_{t=1}^{J-1}u_{st}^{\mathbf{x}}\left(\mathbf{h}_{c}^{\mathbf{x}}\right)^{T}\mathbf{E}_{sJ}\mathbf{h}_{s}^{\mathbf{x}}\ .$
Example 5.
Multinomial logistic models (MLM):To calculate the first-order derivative of $d({\mathbf{x}},\boldsymbol{\xi})$ under MLMs, we follow Appendix C in Huang et al. (2024) after (i) ${\mathbf{F}}(\boldsymbol{\xi})$ is replaced with $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$; (ii) ${\mathbf{F}}_{\mathbf{x}}$ is replaced with $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$; (iii) ${\mathbf{U}}_{\mathbf{x}}=(u_{st}^{\mathbf{x}})_{s,t=1,\ldots,J}$ with $u_{st}^{\mathbf{x}}$ replaced by $E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}$ or $\hat{E}\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}$; and (iv) when $k\geq 1$, for $i=1,\ldots,k$, $\frac{\partial{\mathbf{U}}_{\mathbf{x}}}{\partial x_{i}}=(\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}})_{s,t=1,\ldots,J}$ with $\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}$ replaced by $\frac{\partial E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\int_{\boldsymbol{\Theta}}u_{st}^{\mathbf{x}}(\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ or $\hat{E}\{\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}(\boldsymbol{\Theta})\}=\frac{1}{B}\sum_{j=1}^{B}\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}(\hat{\boldsymbol{\theta}}_{j})$ Note that $\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}(\hat{\boldsymbol{\theta}}_{j})$ denotes $\frac{\partial u_{st}^{\mathbf{x}}}{\partial x_{i}}$ at $\boldsymbol{\theta}=\hat{\boldsymbol{\theta}}_{j}$ , whose formulae can be found in Appendix C of Huang et al. (2024). $\Box$
Example 6.
Generalized linear models (GLM):Under GLMs, the sensitivity function $d({\mathbf{x}},\boldsymbol{\xi})=E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}\left[E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}\right]^{-1}{\mathbf{h}}({\mathbf{x}})$ for $f_{\rm EW}(\boldsymbol{\xi})$ or $\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}\left[\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}\right]^{-1}{\mathbf{h}}({\mathbf{x}})$ for $f_{\rm SEW}(\boldsymbol{\xi})$. Following Section S.5 of the Supplementary Material of Huang et al. (2024), we denote ${\mathbf{A}}_{E}=[{\mathbf{X}}_{{\boldsymbol{\xi}}}^{T}E\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{{\boldsymbol{\xi}}}]^{-1}$ and ${\mathbf{A}}_{\hat{E}}=[{\mathbf{X}}_{{\boldsymbol{\xi}}}^{T}\hat{E}\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}{\mathbf{X}}_{{\boldsymbol{\xi}}}]^{-1}$. The first-order derivative of $d({\mathbf{x}},\boldsymbol{\xi})$ with respect to continuous factors ${\mathbf{x}}_{(1)}=(x_{1},\ldots,x_{k})^{T}$ is $\frac{\partial d({\mathbf{x}},\boldsymbol{\xi})}{\partial{\mathbf{x}}_{(1)}}={\mathbf{h}}({\mathbf{x}})^{T}{\mathbf{A}}_{E}{\mathbf{h}}({\mathbf{x}})\cdot\frac{\partial E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}{\boldsymbol{\Theta}}\}]}{\partial{\mathbf{x}}_{(1)}}+2\cdot E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}{\boldsymbol{\Theta}}\}]\cdot\{\frac{\partial{\mathbf{h}}({\mathbf{x}})}{\partial{\mathbf{x}}_{(1)}^{T}}\}^{T}{\mathbf{A}}_{E}{\mathbf{h}}({\mathbf{x}})$ for $f_{\rm EW}(\boldsymbol{\xi})$, or
$\frac{\partial d({\mathbf{x}},\boldsymbol{\xi})}{\partial{\mathbf{x}}_{(1)}}$ $=$ ${\mathbf{h}}({\mathbf{x}})^{T}{\mathbf{A}}_{\hat{E}}{\mathbf{h}}({\mathbf{x}})\cdot\left\{\frac{\partial{\mathbf{h}}({\mathbf{x}})}{\partial{\mathbf{x}}_{(1)}^{T}}\right\}^{T}\cdot\frac{1}{B}\sum_{j=1}^{B}\nu^{\prime}\{{\mathbf{h}}({\mathbf{x}})^{T}{\hat{\boldsymbol{\theta}}}_{j}\}\hat{\boldsymbol{\theta}}_{j}$
for $f_{\rm SEW}(\boldsymbol{\xi})$. $\Box$
Theorem 6.
For a GLM model under Assumptions (A1) and (B3), consider $\boldsymbol{\xi}\in\boldsymbol{\Xi}$ satisfying $|\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}|>0$, where ${\mathbf{W}}_{{\boldsymbol{\xi}}}$ is either $E\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}$ for integral-based EW D-optimality or $\hat{E}\{{\mathbf{W}}_{{\boldsymbol{\xi}}}(\boldsymbol{\Theta})\}$ for sample-based EW D-optimality (see Example 4). Then $\boldsymbol{\xi}$ is EW D-optimal if and only if $\max_{\mathbf{x}\in\mathcal{X}}E[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}[\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}]^{-1}{\mathbf{h}}({\mathbf{x}})\leq p$ for integral-based D-optimality or $\max_{\mathbf{x}\in\mathcal{X}}\hat{E}[\nu\{{\mathbf{h}}({\mathbf{x}})^{T}\boldsymbol{\Theta}\}]\cdot{\mathbf{h}}({\mathbf{x}})^{T}[\mathbf{X}_{\boldsymbol{\xi}}^{T}{\mathbf{W}}_{{\boldsymbol{\xi}}}\mathbf{X}_{\boldsymbol{\xi}}]^{-1}{\mathbf{h}}({\mathbf{x}})\leq p$ for sample-based EW D-optimality.